OpenStudy (anonymous):

(sin^2x -4sinx+4 )/sinx-2

OpenStudy (anonymous):

Simplify

OpenStudy (aum):

\[ \frac{\sin^2(x) - 4\sin(x) + 4}{\sin(x) - 2} \]Let a = sin(x) \[ \frac{\sin^2(x) - 4\sin(x) + 4}{\sin(x) - 2} = \frac{a^2 - 4a + 4}{a-2} = \frac{(a-2)^2}{a-2} = a-2 = \sin(x) - 2 \]

OpenStudy (anonymous):

Thank you so much.

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

Can I interpret others problems like this as well. By replacing with a variable